Let X be a random variable with pdf

𝑓(π‘₯) = {

2(1 + π‘₯)

27 𝑖𝑓 2 ≀ π‘₯ ≀ 5

0 π‘œπ‘‘β„Žπ‘’π‘Ÿπ‘€π‘–π‘ π‘’

Find (i) 𝑃(𝑋 < 4) (ii) 𝑃(3 < π‘₯ ≀ 4)

The density function of sheer strength of spot welds is given by

𝑓(π‘₯) = {

π‘₯/ 160000 π‘“π‘œπ‘Ÿ 0 ≀ π‘₯ ≀ 400

800 βˆ’ π‘₯/ 160000 π‘“π‘œπ‘Ÿ 400 ≀ π‘₯ ≀ 800

Find the number a such that 𝑃(𝑋 < π‘Ž) = 0.50


1
Expert's answer
2022-03-28T10:07:51-0400

"i:\\P\left( X<4 \right) =\int_4^5{f\left( x \right) dx}=\int_4^5{\frac{2\left( 1+x \right)}{27}dx}=\frac{\left( 1+x \right) ^2}{27}|_{4}^{5}=\frac{11}{27}\\ii:\\P\left( 3<X\leqslant 4 \right) =\int_3^4{\frac{2\left( 1+x \right)}{27}dx}=\frac{\left( 1+x \right) ^2}{27}|_{3}^{4}=\frac{1}{3}\\\\a\leqslant 400:\\P\left( X<a \right) =\int_0^a{f\left( x \right) dx}=\int_0^a{\frac{x}{160000}dx}=\frac{a^2}{320000}\\\frac{a^2}{320000}=0.5\Rightarrow a^2=160000\Rightarrow a=400\\a=400 satisfies\,\,P\left( X<a \right) =0.5"